# Bridging mid and near infrared by combining optomechanics and self mixing

### v1

@@ Abstract
This work describes a self-mixing-assisted optomechanical platform for transferring information between near- and mid-infrared radiation. In particular, the self-mixing signal of a mid-infrared quantum cascade laser is used to detect the oscillation of a membrane induced by radiation pressure exerted by a near-infrared excitation beam, which is amplitude-modulated at the membrane resonance frequency. This technique is, in principle, wavelength-independent and, therefore, can link different spectral regions from both the excitation and probe sides. This versatility can pave the way for future applications of this self-mixing-assisted optomechanical platform in communication and advanced sensing systems.
@@ 1 Introduction
Quantum cascade lasers (QCL) 5 are particularly well-suited to be used in SM setups. Similarly to bipolar semiconductor lasers, they are highly sensitive to optical feedback. On the other hand, since the laser transition takes place between the sublevels of the conduction band created by the semiconductor layers heterostructure, the laser transition lifetime is very short ( ps) 6. For this reason, QCLs can be modulated at high speed (GHz and above) 7, 8 and, in principle, are also able to detect SM signals within the same bandwidth. As a consequence, QCLs have been successfully employed for realizing highly integrated sensitive sensors, as feedback-induced variations in voltage or output power can carry precise displacement and/or optical path information 9, 10.
@@
Recent studies have explored similar configurations involving QCL self-mixing with suspended membranes, demonstrating the feasibility of hybrid photonic-mechanical platforms 10, 9. However, these previous works primarily investigated photothermal effects, which are inherently wavelength-dependent. In contrast, this work investigates the role of radiation pressure, exerted by an excitation beam, as a trigger mechanism for membrane oscillation in an SM-assisted optomechanical platform.
In detail, the SM signal generated with a mid-infrared (mid-IR) QCL is used to detect the optomechanical response (i.e., oscillation) of a trampoline membrane when excited by the radiation pressure of an amplitude-modulated (AM) near-infrared (near-IR) beam. In other words, we test and demonstrate the ability to encode the membrane oscillation signal, induced by the near-IR excitation beam, in the mid-IR QCL via the SM coupling.
This means that the membrane, combined with SM, actively enables the transfer of information, i.e., a communication, between the two beams at different infrared wavelengths.
Moreover, by demonstrating that the radiation pressure exerted by the excitation beam is the phenomenon responsible for the membrane oscillation, we prove that this SM-assisted optomechanical platform benefits from being, in principle, excitation-wavelength independent.
On one hand, this platform can be used as a communication gate between the mid-IR and any other spectral range. On the other hand, the universal nature of the SM effect guarantees a similar wavelength independence for the probe source. This endows the system with pump-probe reciprocity, further expanding the applicability of the proposed platform across most of the frequency spectral range.
@@ 2 Experimental Setup and Methodology Description
In particular, the trampoline membrane is a silicon nitride (Si3N4) membrane with a metal coating (Cr/Au 5-30 nm coating). It is fixed on a pass-through-holed piezoelectric actuator (PZT). The PZT can actuate the membrane oscillation, e.g., for calibration purposes. The membrane is housed in a room-temperature vacuum chamber featuring two selected windows that maintain transparency while enabling simultaneous transmission of the impinging laser beams (a 1-mm-thick cyclic-olefin copolymer optical window is used on the near-IR excitation beam side and a CaF2 window on the mid-IR radiation side). The vacuum chamber is pumped at mbar. More details about the membrane design and fabrication are available in ref. 10. For alignment purposes, the membrane platform (membrane, vacuum chamber, PZT; see Fig. 1) is mounted on a 3-axis stage.
@@
On the other side (right side in Fig. 1), the membrane is illuminated by the MIR QCL probe beam. The QCL is operated at room temperature () in a single-mode regime at a bias current of . The laser mounting arrangement allows us to conveniently exploit the emission from both laser facets. The front-facet-emitted beam is used to probe the membrane oscillation, receiving back the resulting optical feedback. To this extent, this radiation is focused on the membrane via a 50-mm lens (L1), and a germanium wedged window (W) is used in the probe path to prevent any injection of the near-IR radiation into the QCL. Instead, the radiation emitted by the back facet of the QCL can be used for free-space transmission of the MIR radiation carrying the SM signal. In our setup, after a free-space propagation, the SM signal is retrieved by detecting the back-facet-emitted radiation via a fast commercial photovoltaic detector (PVI-4TE-5-2x2 by Vigo System). The detector output signal is used as input signal of a lock-in amplifier. Depending on the application, we highlight that the direct detection of the SM signal from the laser voltage variation could be more convenient, as this configuration requires no extra detector. In particular, the presented optomechanical platform can be conveniently exploited in both configurations, whether we are interested in the SM laser emission or voltage modulation 10, and according to the specific requirements and applications.
As depicted in Fig. 1, our system can be operated in two different configurations. In the first configuration (Conf. 1), the PZT induces the membrane oscillation. This configuration is used for system calibration, as it allows us to quantify the photothermal effect on the membrane of both pump and probe under quasi-continuous-wave operation (see Sec. 3). In the second configuration (Conf. 2), the membrane oscillation is induced by AM modulating the excitation beam (without PZT actuation), as discussed in Section 3. In both cases, the lock-in amplifier is used: 1) to control the modulation signal and its frequency sweep, which is sent to the PZT in Conf. 1 or to the AOM in Conf. 2; 2) to demodulate the self-mixing signal measured via the photodetector. This allows the retrieval of the membrane spectral response (e.g., as shown in Fig. 2a).
In both configurations, the impinging power onto the photovoltaic detector has been kept constant, and the detector is operated in the linear responsivity regime as in refs. 17, 18.
@@ 3 Results and Discussion
After this preliminary characterization, we test the membrane response when illuminated with both the mid-IR probe beam and near-IR excitation beam (Fig. 3). For this test, the membrane oscillation is driven by the PZT controlled via the lock-in amplifier with a constant-in-amplitude frequency-sweeping signal (Conf. 1 in Fig. 1). The mid-IR light impinging on the membrane is fixed at its maximum, i.e. , while the excitation beam power is tuned by changing the amplitude of the AOM RF modulation via the signal generator (see Fig. 1). At this probe power level, the feedback signal is strong, and, at the same time, the probe radiation does not saturate the thermal effects on the membrane, as demonstrated by the redshift of the resonance frequency induced by the near-IR radiation (Fig. 3a). Therefore, an independent analysis of the near-IR laser effect on the membrane resonance is possible. As clearly visible in Fig. 3a, the resonance peak is redshifted in frequency when the near-IR laser power increases. Again, for each value of the near-IR impinging power, we estimate the membrane resonance frequency via a Lorentzian fit of the experimental data. The results are shown in Fig. 3b. The resonance frequency (blue points) linearly decreases as the near-IR power increases at a rate of Hz/mW, as estimated via a linear fit (red line). We remark that the difference between this rate value and the one related to the QCL power variation (Fig. 2) can be explained by different factors: mode matching; the direction of the incident beams on the membrane; the use of different windows in the two paths (transmissivity of CaF2 windows at : 95%, transmissivity of COC window at : 85%); and the fact that, on the 1064 nm side, absorption is also influenced by the presence of a 3-nm Cr layer between the SiN and the 50-nm Au layer 10.
@@
Finally, we test the membrane response in Conf. 2. Specifically, in this configuration, the AM-modulated near-IR beam induces the membrane oscillation, which is probed via the mid-IR beam kept at constant working conditions with a fixed optical power of (Fig. 4). As described in section 2, the AM modulation of the near-IR light is obtained by externally driving the AOM via the lock-in amplifier. Moreover, to reconstruct the resonance peak, the AM modulation is swept in the frequency range of the membrane resonance. The procedure is repeated for different values of AM modulation depth. This allows us to monitor the frequency shift of the membrane resonance peak for different values of the near-IR excitation beam power. As shown in Figs. 4a and 4b, a blue-shift is visible in the peak frequency when increasing the modulation depth from 10% to 50%, corresponding to an average near-IR power from to , respectively. In particular, we can see that, in this case, the frequency shift follows a linear trend when plotted against the impinging near-IR power with a shift rate of Hz/mW. In this second configuration, the membrane oscillation, i.e., the presence of the resonance peak in Fig. 4, is due to radiation pressure exerted by the near-IR beam, while the frequency peak redshift is due to the thermal effect. Indeed, these two phenomena are characterized by two different time scales 10. To prove this, we operate the setup in Conf. 1. In detail, while impinging onto the membrane with both the excitation and probe beams, we oscillate the membrane using the PZT driven with the lock-in amplifier, and we externally apply a slow AM modulation (as reported in the legend of Fig. 5) to the excitation beam by driving the AOM with the signal generator. If the AM modulation is slow enough (i.e., comparable to the thermal relaxation time scale), two peaks are appreciable in the spectrum, as shown in Fig. 5. The first one is at the resonance frequency of the membrane when illuminated with just the probe radiation (no excitation beam). In contrast, the second peak, which is redshifted compared to the previous one, is due to the thermal effect of illuminating the membrane with both the excitation and probe radiation. The net effect of AM modulating the excitation beam is a frequency modulation (FM) of the resonance peak. The spectra are, in fact, characterized by the typical shape of a frequency-modulated signal 19. The spectrum is made of two clearly separated peaks up to , while starting from , and above, they merge into one single peak. This means that the thermal relaxation effect bandwidth is lower than .
Therefore, the presence of the resonance peak when the system is driven in Conf. 2 is only due to the radiation pressure effect, as the timescale is too fast for the thermal effect (more than three orders of magnitude of difference). Instead, the redshift measured both in Conf. 1 and Conf. 2 is due to thermal effects. A more detailed theoretical description of thermal effects on the membrane is available in ref. 10.
To summarize, by AM modulating the excitation radiation in the range of the membrane resonance frequency, we demonstrate the capability of encoding the AM modulation information into the SM signal in the QCL light when the membrane is excited via radiation pressure. Therefore, the presented optomechanical setup promises to serve as an information transducer between the two different wavelengths.
@@ 4 Conclusion and Perspectives
This allows us to monitor, via the SM signal, the membrane oscillations induced by radiation pressure exerted by the near-IR beam when AM-modulated around the membrane resonance frequency. Therefore, by demonstrating the capability of detecting the resonance signal, we show the potential of using this optomechanical system as a communication gate (i.e., information transducer) between two different spectral regions. Indeed, the oscillation of the membrane can be used to amplify and transmit the AM modulation signal of the excitation beam to the probe.
The tested membrane interface is wavelength-independent, therefore allowing a connection between the mid-IR and the excitation beam, regardless of its color.
Moreover, in principle, this system can also be exploited to control several QCL parameters, like frequency, intensity, phase, or power, by SM, which boasts a bandwidth close to 100 kHz.
@@
From a more generic perspective, by engineering optimized membranes for communication purposes (e.g., high resonance frequencies), the techniques presented here could be exploited to create optomechanical interfaces between the fiber-based telecom communication infrastructure and mid-IR free-space communication channels. The latter are already used for point-to-point communication 20, overcoming also adverse weather conditions due to the much reduced scattering 19, 21, 20, 22.
Additionally, starting from this optomechanical platform, novel sensing and imaging setups can be developed: a membrane-based array, for instance, could allow spatial control of the modulation amplitude in the probe, either exploiting the pump intensity distribution or the different mechanical frequencies of each membrane 23. Furthermore, this system can pave the way to advanced imaging methods where the mid-IR laser SM signal is used to probe, e.g, near-IR or visible photothermally excited substrates or targets, envisioning hybrid photoacoustic mid-IR (or THz) sensing 24.
@@
The data that support the findings of this study are available from the corresponding author upon reasonable request.
The author would like to thank Irene La Penna (LENS) for her help in pre-aligning the laser from the laser output to the AOM and Davide Mazzotti (CNR-INO) for the fruitful conversation about possible perspectives.
The authors acknowledge financial support from the European Union—Next Generation EU with the Italian National Recovery and Resilience Plan (NRRP), Mission 4, Component 2, Investment 1.3, CUP D43C22003080001, partnership on “Telecommunications of the Future” (PE00000001 — program “RESTART”), and the I-PHOQS Infrastructure ”Integrated infrastructure initiative in Photonic and Quantum Sciences” [IR0000016, ID D2B8D520], with the Laserlab-Europe Project [G.A. n.871124], with the MUQUABIS Project “Multiscale quantum bio-imaging and spectroscopy” [G.A. n.101070546], with the QUID project "Quantum Italy Deployment" [G.A. No 101091408]; from the Italian ESFRI Roadmap (Extreme Light Infrastructure - ELI Project); from ASI and CNR under the Joint Project “Laboratori congiunti ASI-CNR nel settore delle Quantum Technologies (QASINO)” (Accordo Attuativo n. 2023-47-HH.0); from the Italian Ministero dell’Università e della Ricerca (project PRIN-2022KH2KMT QUAQK).

### v2

@@ Abstract
This work describes a self-mixing-assisted optomechanical platform for transferring information between near- and mid-infrared radiation. In particular, the self-mixing signal of a mid-infrared quantum cascade laser is used to detect the oscillation of a membrane driven by light-induced forces exerted by a near-infrared excitation beam, which is amplitude-modulated at the membrane resonance frequency. This technique benefits from spectral broadness and, therefore, can link different spectral regions from both the excitation and probe sides. This versatility can pave the way for future applications of this self-mixing-assisted optomechanical platform in communication and advanced sensing systems.
@@ 1 Introduction
Quantum cascade lasers (QCL) 5 are particularly well-suited to be used in SM setups. Similarly to bipolar semiconductor lasers, they are highly sensitive to optical feedback. On the other hand, since the laser transition takes place between the sublevels of the conduction band created by the semiconductor layers’ heterostructure, the laser transition lifetime is very short ( ps) 6. For this reason, QCLs can be modulated at high speed (GHz and above) 7, 8 and, in principle, are also able to detect SM signals within the same bandwidth. As a consequence, QCLs have been successfully employed for realizing highly integrated sensitive sensors, as feedback-induced variations in voltage or output power can carry precise displacement and/or optical path information 9, 10.
@@
Recent studies have explored similar configurations involving QCL self-mixing with suspended membranes, demonstrating the feasibility of hybrid photonic-mechanical platforms 10, 9.
However, these previous studies primarily focused on stationary photothermal effects in systems where membrane motion was piezoelectrically driven and only perturbed by external radiation. In contrast, here we explore a different regime in which dynamic light-induced forces—originating from both radiation pressure and photothermal effects 17—serve as the primary driving mechanism for membrane oscillation. Specifically, we demonstrate a fully optical actuation scheme in which a trampoline membrane is driven by amplitude modulation of a near-infrared (near-IR) excitation laser and monitored via self-mixing interferometry using a mid-infrared (mid-IR) QCL.
In other words, we test and demonstrate the ability to encode the membrane oscillation signal, induced by the near-IR excitation beam, in the mid-IR QCL via the SM coupling.
This means that the membrane, combined with SM, actively enables the transfer of information, i.e., a communication between the two beams at different infrared wavelengths.
Moreover, by demonstrating that the light-induced force exerted by the excitation beam is the phenomenon responsible for the membrane oscillation, we prove that this SM-assisted optomechanical platform benefits from being, in principle, broadband in terms of excitation wavelength.
On one hand, this platform can be used as a communication gate between the mid-IR and other spectral ranges. On the other hand, the universal nature of the SM effect guarantees wavelength independence for the probe source. This further expands the applicability of the proposed platform across most of the frequency spectral range.
@@ 2 Experimental Setup and Methodology Description
In particular, the trampoline membrane is a silicon nitride (Si3N4) membrane with a metal coating (Cr/Au 5-30 nm coating). More details about the membrane design and fabrication are available in ref. 10. The membrane is fixed on a pass-through-holed piezoelectric actuator (PZT). The PZT can actuate the membrane oscillation, e.g., for calibration purposes. The membrane is housed in a room-temperature vacuum chamber featuring two selected windows that maintain transparency while enabling simultaneous transmission of the impinging laser beams. In detail, a 1-mm-thick cyclic-olefin copolymer optical window is used on the near-IR excitation beam side and a CaF2 window on the mid-IR radiation side. The vacuum chamber is pumped at mbar. For alignment purposes, the membrane platform (membrane, vacuum chamber, PZT; see Fig. 1) is mounted on a 3-axis stage.
@@
On the other side (right side in Fig. 1), the membrane is illuminated by the MIR QCL probe beam. The QCL is operated at room temperature () in a single-mode regime at a bias current of . The laser mounting arrangement allows us to conveniently exploit the emission from both laser facets. The front-facet-emitted beam is used to probe the membrane oscillation, receiving back the resulting optical feedback. To this extent, this radiation is focused on the membrane via a 50-mm lens (L1), and a germanium wedged window (W) is used in the probe path to prevent any injection of the near-IR radiation into the QCL. Instead, the radiation emitted by the back facet of the QCL can be used for free-space transmission of the MIR radiation carrying the SM signal. In our setup, after a free-space propagation, the SM signal is retrieved by detecting the back-facet-emitted radiation via a fast commercial photovoltaic detector (PVI-4TE-5-2x2 by Vigo System). The detector output signal is used as input signal of a lock-in amplifier. Depending on the application, we highlight that the direct detection of the SM signal from the laser voltage variation could be more convenient, as this configuration requires no extra detector. In particular, the presented optomechanical platform can be conveniently exploited in both configurations 10.
As depicted in Fig. 1, our system can be operated in two different configurations. In the first configuration (Conf. 1), the PZT induces the membrane oscillation. This configuration is used for system calibration, as it allows us to quantify the photothermal effect on the membrane of both pump and probe under quasi-continuous-wave operation (see Sec. 3). In the second configuration (Conf. 2), the membrane oscillation is induced by AM modulating the excitation beam (without PZT actuation), as discussed in Section 3. In both cases, the lock-in amplifier is used: 1) to control the sinusoidal modulation signal and its frequency sweep, which is sent to the PZT in Conf. 1 or to the AOM in Conf. 2; 2) to demodulate the self-mixing signal measured via the photodetector. This allows the retrieval of the membrane spectral response (e.g., as shown in Fig. 2a).
In both configurations, the impinging power onto the photovoltaic detector has been kept constant, and the detector is operated in the linear responsivity regime as in refs. 18, 19.
@@ 3 Results and Discussion
After this preliminary characterization, we test the membrane response when illuminated with both the mid-IR probe beam and near-IR excitation beam (Fig. 3). For this test, the membrane oscillation is driven by the PZT controlled via the lock-in amplifier (Conf. 1 in Fig. 1). The mid-IR light impinging on the membrane is fixed at its maximum, i.e. while the excitation beam power is tuned by changing the amplitude of the AOM RF modulation via the signal generator (see Fig. 1). At this probe power level, the feedback signal is strong, and, at the same time, the probe radiation does not saturate the thermal effects on the membrane, as demonstrated by the redshift of the resonance frequency induced by the near-IR radiation (Fig. 3a). Therefore, an independent analysis of the near-IR laser effect on the membrane resonance is possible. As clearly visible in Fig. 3a, the resonance peak is redshifted in frequency when the near-IR laser power increases. Again, for each value of the near-IR impinging power, we estimate the membrane resonance frequency via a Lorentzian fit of the experimental data. The results are shown in Fig. 3b. The resonance frequency (blue points) linearly decreases as the near-IR power increases at a rate of Hz/mW, as estimated via a linear fit (red line). We remark that the difference between this rate value and the one related to the QCL power variation (Fig. 2) can be explained by different factors: mode matching; the direction of the incident beams on the membrane; the use of different windows in the two paths (transmissivity of CaF2 windows at : 95%, transmissivity of COC window at : 85%); and the fact that, on the 1064 nm side, absorption is also influenced by the presence of a 3-nm Cr layer between the SiN and the 50-nm Au layer 10.
@@
Finally, we test the membrane response in Conf. 2. Specifically, in this configuration, the AM-modulated near-IR beam induces the membrane oscillation, which is probed via the mid-IR beam kept at constant working conditions with a fixed optical power of (Fig. 4). As described in section 2, the AM modulation of the near-IR light is obtained by externally driving the AOM via the lock-in amplifier. Moreover, to reconstruct the resonance peak, the AM modulation is swept in the frequency range of the membrane resonance. The procedure is repeated for different values of AM modulation amplitude. This allows us to monitor the frequency shift of the membrane resonance peak for different values of the near-IR excitation beam power. As shown in Figs. 4a and 4b, a blue-shift is visible in the peak frequency when increasing the AM modulation amplitude from to , corresponding to an average near-IR power from to , respectively. In particular, we can see that, in this case, the frequency shift follows a linear trend when plotted against the impinging near-IR power with a shift rate of Hz/mW. In this second configuration, the membrane oscillation, i.e., the presence of the resonance peak in Fig. 4, is due to light-induced force exerted by the near-IR beam. Moreover, the increase in impinging power leads to a redshift of the resonance frequency, as already shown also in Conf. 1. In general, the light-induced force might have different contributions, i.e., radiation pressure and photothermal force 17. Typically, these two phenomena have different time scales, i.e., photothermal force is characterized by a slower time scale compared to the quasi-instantaneousness of the radiation pressure 17. However, even if their time scale can be quite different depending on the specific oscillating system (e.g., membrane and substrate materials), the thermal effect can be non-negligible—even if dumped— at the resonance frequency10, 17.
Therefore, depending on the applications and specific setup, the driving optomechanical force of the membrane oscillation can be dominated by radiation pressure, photothermal force, or a combination of the two. In view of further investigation to understand the photo-thermal contribution in the system, we report a preliminary characterization of its time scale. The SM setup is set in Conf. 1, i.e. the membrane is oscillated using the PZT driven with the lock-in amplifier while impinging onto the membrane with both the excitation and probe beams. Moreover, we externally apply a slow AM modulation (as reported in the legend of Fig. 5) to the excitation beam by driving the AOM with the signal generator. If the AM modulation is slow enough (i.e., comparable to the thermal relaxation time scale), two peaks are appreciable in the spectrum, as shown in Fig. 5. The first one is at the resonance frequency of the membrane when illuminated with just the probe radiation (no excitation beam). In contrast, the second peak, which is redshifted compared to the previous one, is due to the thermal effect of illuminating the membrane with both the excitation and probe radiation. The net effect of AM modulating the excitation beam is a frequency modulation (FM) of the resonance peak. The spectra are, in fact, characterized by the typical shape of a frequency-modulated signal 20. The spectrum is made of two clearly separated peaks up to , while starting from , and above, they merge into one single peak. This implies that the thermal relaxation effect bandwidth is lower than and significantly below the membrane resonance frequency (). This suggests a significant dump of thermal effect in the time scale characterizing the membrane oscillation; however, deeper investigations are needed to understand which is the main driving force. To summarize, by AM modulating the excitation radiation in the range of the membrane resonance frequency, we demonstrate the capability of encoding the AM modulation information into the SM signal in the QCL light when the membrane is excited via light-induced force. Therefore, the presented optomechanical setup promises to serve as an information transducer between the two different wavelengths.
@@ 4 Conclusion and Perspectives
This enables the monitoring of membrane oscillations fully optically induced by the near-IR beam through the SM signal when the beam is amplitude-modulated at frequencies close to the membrane resonance. Therefore, by demonstrating the capability of detecting the resonance signal, we show the potential of using this optomechanical system as a communication gate (i.e., information transducer) between two different spectral regions. Indeed, the oscillation of the membrane can be used to amplify and transmit the AM modulation signal of the excitation beam to the probe.
The tested membrane interface is wavelength-independent, therefore allowing a connection between the mid-IR and the excitation beam, regardless of its color. Moreover, in principle, this system can also be exploited to control several QCL parameters, like frequency, intensity, phase, or power, by SM, which boasts a bandwidth close to 100 kHz.
@@
From a more generic perspective, by engineering optimized membranes for communication purposes (e.g., high resonance frequencies), the techniques presented here could be exploited to create optomechanical interfaces between the fiber-based telecom communication infrastructure and mid-IR free-space communication channels. The latter are already used for point-to-point communication, also overcoming adverse weather conditions due to the much reduced scattering20, 21, 22, 23. Additionally, starting from this optomechanical platform, novel sensing and imaging setups can be developed: a membrane-based array, for instance, could allow spatial control of the modulation amplitude in the probe, either exploiting the pump intensity distribution or the different mechanical frequencies of each membrane 24. Furthermore, this system can pave the way to advanced imaging methods where the mid-IR laser SM signal could be used to probe, e.g, near-IR or visible photothermally excited substrates or targets, envisioning hybrid photoacoustic mid-IR (or THz) sensing 25.
@@
The data that support the findings of this study is available from the corresponding author upon reasonable request.
The author would like to thank Irene La Penna (LENS) for her help in pre-aligning the laser from the laser output to the AOM, Davide Mazzotti (CNR-INO) for the fruitful conversation about possible perspectives, and Mario Siciliani De Cumis (ASI) and Francesco Marino (CNR-INO) for the useful discussion about thermal effects.
The authors acknowledge financial support from the
European Union—Next Generation EU with the Italian National Recovery and Resilience Plan (NRRP), Mission 4, Component 2, Investment 1.3, CUP D43C22003080001, partnership on “Telecommunications of the Future” (PE00000001 — program “RESTART”), and the I-PHOQS Infrastructure ”Integrated infrastructure initiative in Photonic and Quantum Sciences” [IR0000016, ID D2B8D520], with the Laserlab-Europe Project [G.A. n.871124], with the MUQUABIS Project “Multiscale quantum bio-imaging and spectroscopy” [G.A. n.101070546], with the QUID project "Quantum Italy Deployment" [G.A. No 101091408]; from the Italian ESFRI Roadmap (Extreme Light Infrastructure - ELI Project); from ASI and CNR under the Joint Project “Laboratori congiunti ASI-CNR nel settore delle Quantum Technologies (QASINO)” (Accordo Attuativo n. 2023-47-HH.0); from the Italian Ministero dell’Università e della Ricerca (project PRIN-2022KH2KMT QUAQK).